OpenStudy (anonymous):

decompose 6x/(x-4)(x+2) into partial fractions

OpenStudy (anonymous):

\[(−6/x+2)+12/(x+4)\]

OpenStudy (dumbcow):

\[\rightarrow \frac{A}{x-4}+\frac{B}{x+2} = \frac{6x}{(x-4)(x+2)}\] A(x+2) + B(x-4) = 6x A+B = 6 2A-4B = 0 --> A=4 --> B=2

OpenStudy (dumbcow):

@RaphaelFilgueiras, its (x-4) not (x+4)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yes i do with x+ 4!

OpenStudy (anonymous):

@ananaa you're fine with the next step?

OpenStudy (anonymous):

I thought so...

OpenStudy (anonymous):

Do you want to check your result?

OpenStudy (dumbcow):

next step? oh is this linked to a larger problem involving an integral

OpenStudy (anonymous):

how do you check your results?

OpenStudy (anonymous):

Just post your result, so we can help you :)